OpenStudy (anonymous):

Simplify the expression. 7x to the -3power times 6x to the 3power (sorry dont know how to make it look like an actual problem)

OpenStudy (anonymous):

7(x)^(-3) * 6(x)^3 I believe?

OpenStudy (anonymous):

Yeah that makes more sense, thanks

OpenStudy (anonymous):

Okay, you know how to solve x^3-y^3 type of equations?

OpenStudy (anonymous):

yeah, I know that i would multiply 6 and7 which =42 but i dont know what to do with the -3 and 3

OpenStudy (anonymous):

Take x to be cuberoot(7)*(x)^(-1) and y to be cuberoot(6)*x^1

OpenStudy (anonymous):

would it end up being \[\frac{ 42 }{ x^5 }\] ?

OpenStudy (anonymous):

No. the xs get cancelled out.

OpenStudy (anonymous):

The exponents have a property a^m . a^n = a^(m+n) okay? Now, a here is x m is 3, n is -3

OpenStudy (anonymous):

Clear, I believe?

OpenStudy (anonymous):

i think i got it

OpenStudy (anonymous):

So, what really happened was, (7)(6)x^(+3-3) Which is (42)(x)^0 And that is 42*1 which is 42. Clear?